jared is training to run a marathon. Today he ran 14 miles in 2 hours. After running the first 9 miles at a certain speed, he increased his speed by 4 miles per hour for the remaining five miles. Find the speed which Jared was running for the first nine miles.
Answer by richwmiller(17219)

a+b=2,
r*a=9,
(r+4)*b=5
a = 3/2, b = 1/2, r = 6 r+4=10